OpenStudy (anonymous):

14log 14 x how would i expand this?

sam (.sam.):

\[\huge 14\log _{14}x\] is it?

OpenStudy (anonymous):

im not to sure i have a pdf it may alivate things

OpenStudy (anonymous):

number 9

OpenStudy (anonymous):

my teacher makes these hard to read

sam (.sam.):

its \[\huge 14\log _{14}x\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so hang on

OpenStudy (anonymous):

i would first use the power property correct to put the 14 before the log over the x right?

OpenStudy (anonymous):

i get that right?

sam (.sam.):

There's nothing to simplify unless making the base 10

sam (.sam.):

So, use the rule \[\huge \log _{a}b=\frac{\log _{c}b}{\log _{c}a}\] =========================================== \[\huge 14(\log _{14}x)=14(\frac{logx}{\log14})\]

OpenStudy (anonymous):

so change of base?

sam (.sam.):

yep, then \[14\frac{logx}{\log14}=14\log\frac{x}{14}=14(logx-\log14)\]

OpenStudy (anonymous):

okay that makes some sense now on the pdf can you please take a look at number 10?

OpenStudy (campbell_st):

its simply \[\log _{14} x^{14} \]

sam (.sam.):

\[\huge \log _9 x=5\] \[\huge x=9^{5}\]

OpenStudy (anonymous):

you jsut changed it into exponential

sam (.sam.):

yes

sam (.sam.):

\[\huge \log _{a}b=c, b=a ^{c}\]

sam (.sam.):

The first formula that you learn logarithms

sam (.sam.):

when*

OpenStudy (anonymous):

yeah but it says to find the solution

sam (.sam.):

\[x=9^{5}=59049\]

OpenStudy (anonymous):

oh okay that makes sense

OpenStudy (dumbcow):

@.sam loga/logb does not equal log(a/b)

OpenStudy (anonymous):

Guys while you figure out whose right ill be back

sam (.sam.):

@dumbcow i know you will say that XD

OpenStudy (anonymous):

so was he right?